### v1

# End-to-End Predictive Planner for Autonomous Driving with Consistency Models

Abstract
Trajectory prediction and planning are fundamental components for autonomous vehicles to navigate safely and efficiently in dynamic environments. Traditionally, these components have often been treated as separate modules, limiting the ability to perform interactive planning and leading to computational inefficiency in multi-agent scenarios. In this paper, we present a novel unified and data-driven framework that integrates prediction and planning with a single consistency model. Trained on real-world human driving datasets, our consistency model generates samples from high-dimensional, multimodal joint trajectory distributions of the ego and multiple surrounding agents, enabling end-to-end predictive planning. It effectively produces interactive behaviors, such as proactive nudging and yielding to ensure both safe and efficient interactions with other road users. To incorporate additional planning constraints on the ego vehicle, we propose an alternating direction method for multi-objective guidance in online guided sampling. Compared to diffusion models, our consistency model achieves better performance with fewer sampling steps, making it more suitable for real-time deployment. Experimental results on Waymo Open Motion Dataset (WOMD) demonstrate our method’s superiority in trajectory quality, constraint satisfaction, and interactive behavior compared to various existing approaches.
I Introduction
Figure 1: The overview of our method.
To navigate safely and efficiently in a dynamic environment, autonomous vehicles must effectively predict and interact with diverse road participants, including other vehicles, pedestrians, and cyclists. This typically requires a prediction module to anticipate other agents’ future trajectories and a planning module to generate trajectories for the ego vehicle. While prediction modules have widely adopted data-driven approaches to learn from human-driving data, planning modules typically rely on optimization-based methods. These two modules usually operate in a decoupled, alternating manner [31, 10].
This decoupled method introduces fundamental limitations. The interactive behavior generated by the planner is inherently reactive rather than proactive, as the planner responds to other agents without accounting for how they might react to the ego vehicle. This limitation can lead to the computationally expensive theory of mind (TOM) reasoning [53, 12, 17, 18]. In highly interactive scenarios, such as lane merging where proactive nudging is essential, reactive planning may cause the ego vehicle to be stuck [3, 5]. Existing solutions either struggle to scale with the number of agents [25] or compromise optimality due to a restrictive exploration of the solution search space [49].
In addition, modular approaches optimize prediction and planning independently, which often leads to less effective behavior compared to jointly trained, end-to-end frameworks [30]. The alternation between prediction and planning modules also introduces latency and computational inefficiency, making it challenging to meet the demands of real-time operation.
Our key insight is that modeling the joint distribution of future trajectories for both the ego vehicle and surrounding agents can unify prediction and planning into a single, end-to-end framework. In this approach, the ego trajectory, informed by the desired goal location, represents the plan, while the trajectories of surrounding agents serve as predictions. Additional planning constraints can be imposed on the ego trajectory as needed. When trained on large-scale human driving datasets like the Waymo Open Motion Dataset (WOMD) [15], this joint distribution naturally captures the complex interactive behaviors observed in real-world scenarios.
Diffusion models [51, 20] offer a promising method to sample from this high-dimensional, multimodal joint distribution of ego and surrounding agents’ trajectories. They have demonstrated success in generating high-quality images and videos [39, 21] or robot trajectories [38, 27]. Moreover, their ability to model conditional distributions makes them well-suited for trajectory planning applications, as it allows for the incorporation of crucial context for planning and prediction, including trajectory history, map information, ego vehicle’s goal locations, etc. Compared to transformers, diffusion-based method also supports controllable generation to accommodate additional requirements solely at test time through guided sampling, without any additional training.
However, diffusion models typically require many sampling steps to obtain high-quality samples, making it challenging to meet the demands of real-time operations on autonomous vehicles. To overcome this limitation, consistency models [52] are designed to generate data directly from noise in a single step or a small number of steps, significantly improving sampling efficiency without compromising quality.
In this paper, we present an end-to-end predictive planner based on consistency models, designed to unify planning and prediction within a single data-driven framework. Our method samples from the conditional joint distribution of ego and surrounding agents’ future trajectories, given encoded map and trajectory history along with the ego vehicle’s goal configuration. While we use a transformer-based scene encoder [45] in our implementation, our approach is encoder-agnostic and can integrate with any suitable scene encoder. Trained on WOMD [15], our method ensures proactive trajectory planning by capturing complex interactions between agents observed in the real world. Additionally, during online testing, planning constraints are incorporated through guided sampling using our proposed alternating direction method for multi-objective guidance.
Experiments demonstrate that our consistency model-based planner generates effective interactive behaviors with other agents, such as nudging and yielding, through a single inference, eliminating the need for iterative alternation between prediction and planning modules. It outperforms diffusion models by requiring fewer sampling steps while delivering superior trajectory quality. Compared to transformer-based models, it enables controllable generation through online guided sampling that significantly improves planning constraint satisfaction.
II Related Work
For trajectory predictions, data-driven methods have shown promising results when trained on large-scale real-world human driving datasets such as Waymo Open Motion Dataset (WOMD) [15], INTERACTION Dataset [62], Argoverse [8] or nuScenes [7]. Graph Neural Networks have been used to model the semantic information in traffic scenarios for interactive prediction [23]. Generative models like Variational Autoencoders and Diffusion Models have also been utilized for multimodal trajectory prediction tasks [43, 27]. Recently, transformers [56], particularly Motion Transformers (MTR) [45, 46, 34] with encoder-decoder architectures, have achieved state-of-the-art results in trajectory prediction [58].
In trajectory planning, the goal is to determine a feasible and efficient path for the ego vehicle, based on observations of the environment and surrounding agents. This task is often formulated as a numerical optimization problem, typically decomposed path planning and speed planning to facilitate real-time operation in dynamic environments [60, 61, 2]. Game-theoretic approaches have been explored to model agent interactions [41], with efficient computation [16] that considers imperfect traffic agents [54], safe actions [55], and uncertainty reduction in the estimates of agent behaviors [22].
Beyond standalone planning, preliminary effort has been made to integrate trajectory predictions into planner design for collision avoidance [31] and proactive motion planning [4, 44, 49]. While the aforementioned methods have advanced trajectory prediction and planning performance in dense environments, to our best knowledge, none have attempted to unify these two to achieve human-like interactive behavior with computational efficiency. Instead, our work addresses this gap by presenting a data-driven framework for end-to-end joint prediction and planning with a single inference.
Diffusion models [47, 51, 20] are generative models capable of producing high-quality samples from complex distributions, initially achieving success in image and video generations [39, 21]. The Denoising Diffusion Probabilistic Model (DDPM) [20] gradually corrupts the training data with Gaussian noise and learns to reverse this process through iterative denoising with deep neural networks. These models commonly use U-Net [40] or transformer-based architectures [56] as backbones. Conditional sampling in diffusion models is enabled by techniques such as classifier guidance [13] and classifier-free guidance [19] However, DDPM requires numerous sampling steps to preserve the sample quality, resulting in significant computational costs. To address this inefficiency, Denoising Diffusion Implicit Models (DDIM) [48] introduces a deterministic generation process that balances sampling quality and speed, while progressive distillation [42] reduces the number of sampling steps by distilling a pre-trained diffusion model into a more efficient one. Consistency model [52, 50] takes a different approach, directly generating data from noise. They can be trained through distillation of a pre-trained diffusion model or a standalone consistency training and outperform distillation methods in one- and few-step sampling [52].
Diffusion and consistency models also have shown success in robotics and autonomous systems for their expressiveness and ability to provide samples from complex distributions. In robotics, conditional diffusion models have been applied to policy learning [26, 1, 11], with consistency models following similar applications [14, 37]. Diffusion models have also been adopted for trajectory optimization across various robotics tasks [33, 32]. In autonomous driving, diffusion models have shown promising results in trajectory prediction [27, 57] and controllable traffic simulation or scenario generation [63, 59, 9, 24], capturing complex interactive behaviors in traffic scenarios. However, our work differs fundamentally from these efforts: we focus on developing a predictive trajectory planner with consistency models, leveraging guided sampling to incorporate planning constraints. Unlike scenario generation tasks, where sampling efficiency is less critical, our consistency model-based predictive planner needs significantly shorter sampling steps to deliver high-quality interactive behaviors compared to diffusion models.
III Problem Formulation
We aim to plan the trajectory of the ego vehicle with a planning goal, simultaneously predicting the interactive behaviors of others. We assume the trajectory histories of traffic agents and map information are given and known.
At the current timestep $t_{0}\in\mathbb{R}$, the ego agent’s state history over the past $H_{1}\in\mathbb{N}$ timesteps including the current state is given as $S_{e}^{\text{hist}}=(s_{e}^{t_{0}},s_{e}^{t_{0}-1},...,s_{e}^{t_{0}-H_{1}})\in\mathbb{R}^{(H_{1}+1)\times d_{s}}$, where $d_{s}$ is the dimension of the states, e.g. position, velocity, etc. Let $\mathcal{O}$ denote the set of other agents of all types such as vehicles, pedestrians, and cyclists. Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$. The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c$ points on each polyline and $d_{m}$ attributes per point. We use fixed dimensions $N_{\mathcal{O}}$ and $N_{m}$ for agents and map polylines respectively, which are chosen large enough to accommodate all scenarios with padding and masks applied to invalid entries. We assume the ego agent’s goal position is given by some high-level planner and is notated $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
Let $\mathcal{O}_{s}\subset\mathcal{O}$ denote the subset of surrounding agents that may interact with the ego agents. Given the information $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M,s_{e}^{g})$, for the future $H_{2}\in\mathbb{N}$ timesteps, we aim to plan the ego agent’s trajectory $S_{e}^{\text{future}}=(s_{e}^{t_{0}+1},s_{e}^{t_{0}+2},...,s_{e}^{t_{0}+H_{2}})\in\mathbb{R}^{H_{2}\times d_{s}}$, and predict the trajectories of surrounding agents $S_{\mathcal{O}_{s}}^{\text{future}}=(s_{o}^{t_{0}+1},s_{o}^{t_{0}+2},...,s_{o}^{t_{0}+H_{2}})_{o\in\mathcal{O}_{s}}\in\mathbb{R}^{N_{\mathcal{O}_{s}}\times(H_{2})\times d_{s}}$, where $N_{\mathcal{O}_{s}}=|\mathcal{O}_{s}|$. This joint prediction and planning naturally captures interactive behaviors between agents. Additionally, the planned trajectory for the ego vehicle should satisfy $N_{c}\in\mathbb{N}$ constraints $c_{i}(S_{e}^{\text{future}})\leq 0,i=1,2,...,N_{c}$, which may include goal reaching, control limits, or other requirements.
IV Preliminaries
IV-A Motion-Transformer Encoder
To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [45, 46]. This transformer-based architecture effectively models scene context through a local attention mechanism, enabling efficient interaction modeling between agents and the road map while maintaining memory efficiency. It also introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually. For more detailed descriptions of this encoder, we refer readers to the original MTR paper [45]. The MTR++ encoder [46] would be suitable for multi-agent settings, but its implementation is not yet available. More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to integrate with our consistency model-based predictive planner. We represent the encoded feature of the trajectory history and map with MTR encoder as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
IV-B Consistency Model
The consistency model can generate high-quality samples from complex distributions with only one or a few sampling steps [52]. It consists of a forward diffusion process and a reverse process. Let $\mathcal{D}$ be our dataset of trajectories. We first normalize all trajectories in $\mathcal{D}$ using the empirical mean and standard deviation computed across the entire dataset. Let $\mathcal{X}$ be the space of such normalized trajectories, and $p_{\text{data}}$ be the data distribution over $\mathcal{X}$. In the forward process, we first draw an initial sample $x_{1}$ from $p_{\text{data}}$. We then apply an increasing noise schedule ${\sigma_{1},\sigma_{2},...,\sigma_{T}}$ with $\sigma_{1}\approx 0$ to progressively corrupt $x_{1}$ through $T$ steps. Specifically, at each step $i$, we draw a noise sample $\epsilon$ from $\mathcal{N}(0,I)$ and set the corrupted data $x_{i}$ as:
$\quad x_{i}=x_{1}+\sigma_{i}\cdot\epsilon,\quad i=1,2,...,T$ (1)
We choose $\sigma_{T}$ large enough so that, when $x_{1}$ is repeatedly sampled from $p_{\text{data}}$ and corrupted via additive Gaussian noise, the empirical distribution of the resulting $x_{T}$ approximates $\mathcal{N}(0,\sigma_{T}^{2}I)$.
Let $\mathcal{Y}$ be the space of conditioning information. In the reverse process, we aim to learn a consistency function $f_{\theta}$ with the parameter $\theta$ that maps a noisy trajectory sample $x_{i}$, condition $y\in\mathcal{Y}$, and the noise level $\sigma_{i}$ directly to the corresponding clean sample $x_{1}$ for each step $i=1,..,T$. This is achieved by choosing a specific function form of $f_{\theta}$:
$f_{\theta}(x,y,\sigma)=c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(x,y,\sigma)$ (2)
where $\sigma$ represents the noise level, and the differentiable function $c_{\text{skip}}$ and $c_{\text{out}}$ are chosen such that $c_{\text{skip}}(\sigma_{1})=1,c_{\text{out}}(\sigma_{1})=0$ when $\sigma=\sigma_{1}$ [28]. This ensures that the boundary condition of $f_{\theta}$ is met: $f_{\theta}(x_{1},y,\sigma_{1})=x_{1}$. $F_{\theta}$ is usually approximated by the neural network using a U-Net [40] or a transformer [36] based architecture.
For the consistency model training, we aim to enforce the consistency of the output from $f_{\theta}$ for adjacent sampling steps $i$ and $i+1$ to be both $x_{1}$. Let $\mathcal{P}[1,T-1]$ be a distribution over the integer set ${1,2,...,T-1}$. The consistency training minimizes the following loss function $\mathcal{L}_{\text{consistency}}$:
$\mathcal{L}_{\text{consistency}}=$
$\mathbb{E}_{i\sim\mathcal{P}[1,T-1],\epsilon\sim\mathcal{N}(0,I)}\bigl{[}d(f_{\theta}(x_{i},y,\sigma_{i})-f_{\theta}(x_{i+1},y,\sigma_{i+1}))\bigr{]}$ (3)
where $\mathcal{P}$ can be the uniform distribution or lognormal distribution, and $d$ is chosen to be a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [50].
During the data generation process, we first draw a sample $x_{T}$ from $\mathcal{N}(0,\sigma_{T}^{2}I)$. Then with the trained consistency model $f_{\theta}$, we perform iterative sampling by first predicting the approximate clean data $x_{1}^{[i]}$ and then sample $x_{i-1}$ with $\epsilon$ drawn from $\mathcal{N}(0,I)$:
$x_{1}^{[i]}=f_{\theta}(x_{i},y,\sigma_{i}),\quad x_{i-1}=x_{1}^{[i]}+\sigma_{i-1}\epsilon$ (4)
for $i=T,T-1,...,1$, until we obtain the final clean sample $x_{1}^{[1]}$.
V Methodology
With the aforementioned notations and stated problem, our proposed method is summarized in Fig. 1. We first use an MTR encoder to encode the agent’s trajectory history and map information. Then we use a consistency model that takes the conditional input of the ego agent’s planning goal and the MTR-encoded features and generates trajectory plans and predictions for the ego and surrounding agents, respectively. Additional planning constraints for the ego agent are achieved through guided sampling of consistency models.
V-A Data pre-processing
For training diffusion models or consistency models, we usually need to normalize the input data. If we jointly model ego and surrounding agents’ future trajectories $(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$, using an ego-centric coordinate system introduces large variances in the data, particularly for surrounding agents whose positions vary significantly across scenarios. Thus the training performance is largely degraded. Instead, we apply a coordinate transformation $\Gamma$ similar to MTR++ [46] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$. This transformation provides us with data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ that has substantially reduced variance. We then compute the empirical mean and standard deviation of the transformed trajectories across the dataset and standardize them to zero mean and unit variance. To preserve the relative spatial relationships between agents, we collect the reference states consisting of each agent’s position at timestep $t_{0}$ as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$.
V-B Consistency Model Training
Let $\mathcal{X}$ denote the space of transformed future trajectories of the ego vehicle $\Gamma(S_{e}^{\text{future}})$ and surrounding agents $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec. V-A. Let $\mathcal{Y}$ denote the space of conditional inputs, containing encoded historical trajectories $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$, map features $\text{MTR}(M)$ from Sec. IV-A, ego agent’s goal states $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec. V-A.
Given a planning goal and environment context as the condition $y\in\mathcal{Y}$, we leverage the consistency model [52, 50] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$. Each trajectory sample represents one possible joint future behavior of the ego vehicle and surrounding agents.
For training our predictive planner, we jointly train the MTR encoder and the consistency model together in one step with a hybrid loss function, which is constructed as a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq. (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec. IV-A as follows:
$\mathcal{L}=\omega_{0}\mathcal{L}_{\text{consistency}}+\omega_{1}\mathcal{L}_{\text{encoder}}$ (5)
where $\omega_{0}$ and $\omega_{1}$ control the weight of these two losses.
V-C Guided Sampling
When generating trajectories with the trained consistency model as described in Sec. IV-B, to impose planning constraints on the ego vehicle’s future trajectory, we present a novel guided sampling approach similar to classifier guidance [13]. Importantly, this guidance is applied solely during sampling at test time, without requiring any modification to the training procedure.
Assuming there are $N_{c}\in\mathbb{N}$ planning constraint $c_{j}$ to minimize, $j=1,2,...,N_{c}$ and drawing inspiration from classifier guidance [13], we can perform a gradient descent step on the predicted $x_{1}^{[i]}$ during each sampling step $i$ in Eq. (4):
$x_{1}^{[i,1]}=x_{1}^{[i]}-\sum^{N_{c}}_{j=1}\alpha_{j}\nabla c_{j}$ (6)
where $\nabla c_{j}$ represents the first-order gradient of the planning constraint $c_{j}$ and can be computed using auto-differentiation. $\alpha_{j}$ is the corresponding step size for each constraint $c_{j}$. Then the tuned $x_{1}^{[i,1]}$ is used to sample $x_{i-1}$ in Eq. (4).
However, optimizing multiple constraints simultaneously presents a significant challenge in finding a suitable stepsize $\alpha$ efficiently, particularly when constraints may conflict with each other. To address this challenge, inspired by the Alternating Direction Method of Multipliers (ADMM) [6], we propose a novel alternating direction method that only optimizes one constraint at a time during each sampling step $i$:
$x_{1}^{[i,1]}=x_{1}^{[i]}-\alpha_{1}\nabla c_{1}$
$x_{1}^{[i,2]}=x_{1}^{[i,1]}-\alpha_{2}\nabla c_{2}$
$x_{1}^{[i,N_{c}]}=x_{1}^{[i,N_{c}-1]}-\alpha_{N_{c}}\nabla c_{N_{c}}$ (7)
In our approach, we sequentially optimize each constraint $c_{j}$ with corresponding step sizes $\alpha_{j}$ during each gradient descent iteration. While our selection of the optimization order and relatively small step sizes is based on empirical observations, this strategy has demonstrated effective convergence in practice, which is difficult to achieve with the standard guidance method with gradient descent in Eq. (6). The convergence of alternating direction methods, such as ADMM, is generally guaranteed under conditions including closed, proper, and convex functions and appropriately chosen step sizes [6].
V-D Planning Constraints Construction
We assume that the ego vehicle’s dynamics are as follows:
$\dot{p}_{x}=v\cos\theta,\quad\dot{p}_{y}=v\sin\theta,\quad\dot{v}=a,\quad\dot{\theta}=\omega$ (8)
where $p_{x}$ and $p_{y}$ are positions, $v$ is the linear speed and $\theta$ is the orientation. The control input is the acceleration $a$ and angular speed $\omega$.
Suppose we only use $p_{x}$ and $p_{y}$ from ego agent’s future states $S_{e}^{\text{future}}$ to construct our planning constraint function $c_{i}$, since other states like $v$ or $\theta$ may not satisfy the dynamics equations in Eq. (8) and are also quite noisy. Using the property of differential flatness for dynamics in Eq. (8), we can infer $v$ and $\theta$ with $p_{x}$ and $p_{y}$ as follows:
$v=\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}},\quad\theta=\arctan(\frac{\dot{p}_{y}}{\dot{p}_{x}})$ (9)
Then we can further infer the control input $a$ and $\omega$:
$a=\frac{\dot{p}_{x}\ddot{p}_{x}+\dot{p}_{y}\ddot{p}_{y}}{\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}},\quad\omega=\frac{\dot{p}_{x}\ddot{p}_{y}-\dot{p}_{y}\ddot{p}_{x}}{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}$ (10)
Note that, $a$ and $v$ are computed through finite differencing of $p_{x}$ and $p_{y}$ which will inevitably incur some noises.
With $p_{x},p_{y}$ and the derived $a$ and $\omega$, we consider three types of planning constraint $c$ to minimize as follows.
$c_{\text{goal}}=\sqrt{(p_{x}^{t_{0}+H_{2}}-p_{x,goal})^{2}+(p_{y}^{t_{0}+H_{2}}-p_{y,goal})^{2}}$
$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (11)
$c_{\text{goal}}$ represents a goal-reaching constraint such that the ego vehicle needs to reach a pre-specified goal $(p_{x,\text{goal}},p_{y,\text{goal}})$ at the final timestep $t_{0}+H_{2}$. $c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints that compute the average control limit violation along the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$.
VI Experiments
VI-A Experiment setup
VI-A1 Dataset
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [15], which contains 486k traffic scenarios in the training set and 44k scenarios in the validation set that we can test on. Each scenario provides road graph information and trajectories of diverse agents including vehicles, pedestrians, and cyclists, with up to 8 agents specifically marked for trajectory prediction. For each agent, this dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals. This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps, respectively. For each scenario, with the trajectory information from all agents including vehicles, pedestrians, and cyclists as $S_{\mathcal{O}}$, we only select vehicle agents as the ego agent with trajectories denoted by $S_{e}$. We then identify the surrounding agents, with trajectories denoted as $S_{\mathcal{O}_{s}}$, by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold. This enables us to focus our predictive planner on the most relevant interactive behaviors between the ego vehicle and its immediate surroundings. The goal $s_{e}^{g}$ of the ego agent is chosen as its state at the end of the 8-second future horizon.
VI-A2 Model Details
The MTR encoder consists of 6 transformer layers with a dimension of 256 to encode agent trajectory histories and map information into $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M)$. For detailed encoder architecture specifications, we refer readers to [45]. The conditional input $y$ for the consistency model is constructed by further encoding three components: the obtained features $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M)$, the ego agent’s goal $s_{e}^{g}$, and the reference states $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$, using 3-layer MLPs with 256 neurons per layer. To model the consistency model function $F_{\theta}$, we adopt the U-Net [40] with the model dimension of 128 as the backbone.
For the consistency model, we choose $T=5$ which indicates 4 sampling steps in the forward and reverse process. Follow the original paper [50], we set the noise schedule with $\rho=6$, $\sigma_{1}=0.002,\sigma_{T}=80$, and then $\sigma_{i}=\left(\sigma_{0}^{1/\rho}+\frac{i-1}{T-1}(\sigma_{T}^{1/\rho}-\sigma_{0}^{1/\rho})\right)^{\rho},i=1,...,T$. During guided sampling, we perform 100 gradient steps of Eq. (V-C) at each sampling step, and set the coefficient for $c_{\text{goal}},c_{\text{acc}}$ and $c_{\omega}$ to be $2\times 10^{-5},3\times 10^{-6},5\times 10^{-7}$, respectively.
Figure 2: Ego vehicle’s planned trajectory (Green) in a left-turn scenario compared to the Groundtruth (Red).
Figure 3: Planning with different goal locations using consistency models and guided sampling.
TABLE I: Planning Trajectory Compared to the Human Groundtruth
Method minADE $\downarrow$ minFDE $\downarrow$
Transformer 0.944 3.239
DDPM-10 0.346 0.377
DDPM-4 0.363 0.498
DDIM-4 2.126 2.559
Consistency 0.309 0.218
Consistency (guided) 0.303 0.016
(a) A proactive nudging behavior from the ego car to switch a lane.
(b) A yielding behavior for the ego car to make a turn.
Figure 4: For each scenario, upper row: interactive behavior from our predictive planner; lower row: Ground truth driving behaviors. The ego agent’s trajectory is in green annotated by “E” and the surrounding agents’ trajectories are in orange annotated by“S”. The other agents’ trajectories (always shown with ground truth) are in blue.
VI-A3 Methods and Baselines
We compare our consistency model-based approach against several baselines. Currently, we only provide the evaluation of the open-loop trajectory generation performance of those methods, focusing on the behavior through a single inference step.
Transformer
A model based on MTR’s encoder-decoder architecture [45] to generate trajectory only for ego vehicle. It has a larger model size for the decoder (512 vs. our 128), but doesn’t incorporate ego vehicle’s goal $s_{e}^{g}$ as input.
DDPM
This diffusion baseline uses the same MTR encoder and MLP architecture as ours for condition encoding but adopts DDPM [20] for trajectory sampling of ego and surrounding agents. Note that it differs from consistency models in both its forward diffusion process and training loss function. We evaluate two variants: one trained and sampled with 4 steps (DDPM-4), and another trained and sampled with 10 steps (DDPM-10).
DDIM
This baseline extends DDPM-10 by incorporating the more efficient DDIM sampling strategy [48]. While using the same training process as DDPM-10, it employs 4-step DDIM sampling during inference, denoted as DDIM-4.
Consistency
Our proposed consistency model is trained and sampled with 4 steps.
Consistency (guided)
Our proposed consistency model is trained and sampled with 4 steps, with guided sampling.
VI-A4 Training
We conduct distributed training with a total batch size of 100 across 4 NVIDIA L40 GPUs, each with a batch size of 25. For consistency and diffusion models, we use the Adam optimizer [29] with a learning rate of $8\times 10^{-5}$ and train for 50 epochs. For the transformer, we follow the MTR settings [45] and use the AdamW optimizer [35] with a learning rate of $1\times 10^{-4}$ and weight decay of 0.01, and train for 30 epochs.
VI-B Experiment results and analysis
In this section, we evaluate our consistency model-based planner from several aspects. First, we assess its planning performance and trajectory quality compared to several baseline methods. Then we take a closer look at interactive scenarios and present examples of proactive nudging and yielding behaviors of our predictive planner. Furthermore, we show how online guided sampling improves constraint satisfaction in planning - the controllability not provided with transformer-based approaches. Finally, we discuss the computational time during inference for our presented methods.
VI-B1 Trajectory Planning Performance
In Fig. 2, we compare planned trajectories generated by different methods (shown in green) against the ground truth from the dataset (shown in red). The scenario requires the ego vehicle to make a left turn and then drive straight to reach its goal. According to Fig. 2, DDPM-4 reaches the goal but produces a noisy trajectory, likely due to insufficient sampling steps for diffusion models to generate high-quality trajectories. While DDPM-10’s increased sampling steps yields slightly smoother trajectories, it comes at the cost of longer computation time. DDIM-4 attempts to accelerate DDPM-10 using only 4 sampling steps, but fails to generate goal-reaching trajectories. The transformer with a larger model size generates a trajectory close to the groundtruth but misses the exact goal location. In contrast, our consistency model generates smooth trajectories that both align with the ground truth and precisely reach the target location.
In Table I, we quantitatively evaluate the planning performance for the ego vehicle compared to the human groundtruth through metrics from Waymo’s Motion Prediction Challenge [58]: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) on the entire test dataset. Here we view these metrics as an evaluation of how closely our generated trajectories align with human driving behavior under the same environmental context and goal conditions. The results align with our qualitative observations: our consistency model achieves the lowest minADE and minFDE, indicating that it captures human driving patterns well. Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints.
In Table II, we evaluate the trajectory quality using three metrics that characterize driving behavior: 1) angle changes, which measure the smoothness of directional transitions, 2) path length, which quantifies the efficiency of the chosen route, and 3) curvature, which assesses the sharpness of turns. Lower values in these metrics generally indicate smoother and more efficient trajectories that align with natural driving patterns. Our consistency model achieves superior trajectory quality compared to all diffusion-based methods, including DDPM-10 despite its additional sampling steps. while the transformer method has slightly lower angle changes and curvature, this is expected given its substantially larger model size (512 vs. our 128 dimensions).
TABLE II: Planning Trajectory Quality
Method Angle changes $\downarrow$ Path len $\downarrow$ Curvature $\downarrow$
Transformer 0.419 59.621 0.560
DDPM-10 0.660 87.801 3.444
DDPM-4 0.729 79.731 2.211
DDIM-4 0.683 157.78 2.012
Consistency 0.636 58.060 1.174
Consistency (guided) 0.445 57.564 0.610
Furthermore, we demonstrate the flexibility of our approach in choosing goal location in Fig. 3. Our consistency model maintains high-quality trajectory planning when targeting novel goal locations, even when they are not present in the dataset. The current transformer method with MTR doesn’t support the selection of goal locations.
VI-B2 Interactive Behavior
A key advantage of our predictive planner is its ability to ensure safe and effective interactions with other road users in traffic. Our planner demonstrates the capability for generating both proactive behaviors as shown in Fig. 4, such as nudging for lane switching and yielding to pedestrians. The ego agent (labeled “E”) trajectory is in green and the surrounding agents (labeled “S”) trajectories are in orange. Other agent’s trajectories are shown in blue. These learned behaviors are similar to ground truth human drivers and ensure efficient navigation and safe interaction between the ego vehicle and other road users. Crucially, all planning and prediction are achieved through a single inference of the consistency model.
In Fig. 4a, we show a proactive nudging scenario. Here, the ego vehicle needs to merge into the left lane while a nearby vehicle occupies that lane. Our planner directs the ego vehicle to accelerate and nudge into the left lane, predicting that the vehicle in the left lane will decelerate to accommodate the lane change. This nudging behavior demonstrates the proactive planning capability of our end-to-end, data-driven predictive planner, which is challenging to achieve using modular methods. While in the ground truth, the left car doesn’t decelerate that much and the rightmost vehicle eventually turns right, this discrepancy is expected given that we currently generate 8-second future trajectories through a single inference.
In Fig. 4b, a yielding behavior is presented. The ego vehicle intends to make a right turn but predicts that a pedestrian will continue crossing the road. As a result, the planner instructs the ego vehicle to wait until the pedestrian has fully crossed, which is highly aligned with the ground truth. This yielding behavior ensures safe interaction between the ego vehicle and other road users, highlighting the planner’s safe operation in dynamic environments.
VI-B3 Guided Sampling
The guided sampling approach described in Sec. V-C and Sec. V-D allows us to enforce planning constraints during inference time, providing a key advantage over transformer-based methods which lack this capability. We quantitatively evaluate all methods against three planning constraints from Eq. (V-D): goal reaching, acceleration limits, and orientation speed limits. As shown in Table III, our base consistency model already achieves more accurate goal-reaching and smoother behaviors than other diffusion-based approaches. With guided sampling, these planning constraint violations of the consistency model further decrease by a factor of 4, achieving the lowest values among all methods. Note that this improvement requires no model retraining and only applies during inference.
TABLE III: Planning Constraint Violations
Method Goal Reaching $\downarrow$ Acc limit $\downarrow$ $\omega$ limit $\downarrow$
Transformer 9.256 0.570 1.029
DDPM-10 12.669 80.581 3.307
DDPM-4 13.873 56.170 3.323
DDIM-4 22.933 218.823 5.270
Consistency 0.618 5.754 2.050
Consistency (guided) 0.127 0.462 0.599
VI-B4 Computational time discussions
All methods in our evaluation support batch sampling. As shown in Table IV, we evaluate their inference times using a batch of 30 scenarios with 125 ego agents on an NVIDIA GeForce RTX 3060. While the transformer achieves the fastest inference time, it lacks our method’s key capabilities: online controllable generation and sampling from multimodal joint distribution. Our consistency model achieves a similar inference time with DDPM-4 and DDIM-4 due to the same model dimensionality, and sampling steps. However, DDPM-4 and DDIM-4 fail to match our consistency model’s performance in trajectory generation across all metrics, as evidenced in Fig. 2, Table I, Table II, and Table III. The guided sampling, introducing a little computational overhead, provides higher quality trajectory generation and enables controllable planning. Even DDPM-10, despite its increased sampling steps and higher computational time, cannot achieve a comparable performance of trajectory generation to our approach.
TABLE IV: Computational Time Comparison for Inference
Method Inference Time (s)
Transformer 1.910
DDPM-10 3.352
DDPM-4 2.226
DDIM-4 2.118
Consistency 2.301
Consistency (Guided) 2.635
VII Limitations
Although our proposed method significantly improves the efficiency of generating human-like trajectories through consistency models and joint prediction and planning, it may not yet meet the demands of real-time operation, especially in highly dynamic environments that require high-frequency planning. Future work could leverage model pruning or distillation techniques to further improve the inference efficiency. Besides, while in the current paper, we test open-loop trajectory generation to focus on evaluating the model performance of a single inference step, in the future we will extend our method to the closed-loop case with a receding horizon fashion, which is expected to further enhance the interactive planning.
VIII Conclusions
In this paper, we present a novel data-driven approach that unifies trajectory prediction and planning using consistency models. By modeling the joint distribution of ego and surrounding agents’ trajectories, our method enables efficient and interactive planning, such as proactive nudging and yielding, without the need for iterative alternation between prediction and planning modules. Trained on the Waymo Open Motion Dataset, the proposed consistency model generates human-like trajectories while its guided sampling capability naturally incorporates additional planning constraints during online trajectory generations. Experimental results demonstrate that our approach achieves superior trajectory quality and interactive behavior compared to existing methods while requiring fewer sampling steps. This framework also has the potential to be extended to broader robot decision-making tasks in interactive scenarios.
Looking forward, we aim to optimize this method for real-time operation and extend it to closed-loop operation with receding horizon planning. Additionally, we plan to incorporate LLMs as high-level planners to interpret diverse contextual information and enhance decision-making in challenging scenarios.

### v2

@@
-End-to-End Predictive Planner for Autonomous Driving with Consistency Models
+Predictive Planner for Autonomous Driving with Consistency Models
 Abstract
-Trajectory prediction and planning are fundamental components for autonomous vehicles to navigate safely and efficiently in dynamic environments.
-Traditionally, these components have often been treated as separate modules, limiting the ability to perform interactive planning and leading to computational inefficiency in multi-agent scenarios.
-In this paper, we present a novel unified and data-driven framework that integrates prediction and planning with a single consistency model.
-Trained on real-world human driving datasets, our consistency model generates samples from high-dimensional, multimodal joint trajectory distributions of the ego and multiple surrounding agents, enabling end-to-end predictive planning.
-It effectively produces interactive behaviors, such as proactive nudging and yielding to ensure both safe and efficient interactions with other road users.
-To incorporate additional planning constraints on the ego vehicle, we propose an alternating direction method for multi-objective guidance in online guided sampling.
-Compared to diffusion models, our consistency model achieves better performance with fewer sampling steps, making it more suitable for real-time deployment.
-Experimental results on Waymo Open Motion Dataset (WOMD) demonstrate our method’s superiority in trajectory quality, constraint satisfaction, and interactive behavior compared to various existing approaches.
+Trajectory prediction and planning are essential for autonomous vehicles to navigate safely and efficiently in dynamic environments.
+Traditional approaches often treat them separately, limiting the ability for interactive planning.
+While recent diffusion-based generative models have shown promise in multi-agent trajectory generation, their slow sampling is less suitable for high-frequency planning tasks.
+In this paper, we leverage the consistency model to build a predictive planner that samples from a joint distribution of ego and surrounding agents, conditioned on the ego vehicle’s navigational goal.
+Trained on real-world human driving datasets, our consistency model generates higher-quality trajectories with fewer sampling steps than standard diffusion models, making it more suitable for real-time deployment.
+To enforce multiple planning constraints simultaneously on the ego trajectory, a novel online guided sampling approach inspired by the Alternating Direction Method of Multipliers (ADMM) is introduced.
+Evaluated on the Waymo Open Motion Dataset (WOMD), our method enables proactive behavior such as nudging and yielding, and also demonstrates smoother, safer, and more efficient trajectories and satisfaction of multiple constraints under a limited computational budget.
 I Introduction
-Figure 1: The overview of our method.
+Figure 1: Our consistency model takes the map and agents history encoded with the planning goal, and generates a trajectory plan for the ego vehicle and predictions for surroundings.
+Multiple constraints are enforced through online guided sampling.
 To navigate safely and efficiently in a dynamic environment, autonomous vehicles must effectively predict and interact with diverse road participants, including other vehicles, pedestrians, and cyclists.
 This typically requires a prediction module to anticipate other agents’ future trajectories and a planning module to generate trajectories for the ego vehicle.
-While prediction modules have widely adopted data-driven approaches to learn from human-driving data, planning modules typically rely on optimization-based methods.
-These two modules usually operate in a decoupled, alternating manner [31, 10].
-This decoupled method introduces fundamental limitations.
-The interactive behavior generated by the planner is inherently reactive rather than proactive, as the planner responds to other agents without accounting for how they might react to the ego vehicle.
-This limitation can lead to the computationally expensive theory of mind (TOM) reasoning [53, 12, 17, 18].
+Traditionally, these two modules usually operate in a decoupled, alternating manner [27, 9], which introduces fundamental limitations.
+Some approaches integrate data-driven predictions into an optimization-based planner that is inherently reactive rather than proactive [15].
 In highly interactive scenarios, such as lane merging where proactive nudging is essential, reactive planning may cause the ego vehicle to be stuck [4, 5].
-Existing solutions either struggle to scale with the number of agents [25] or compromise optimality due to a restrictive exploration of the solution search space [49].
-In addition, modular approaches optimize prediction and planning independently, which often leads to less effective behavior compared to jointly trained, end-to-end frameworks [30].
-The alternation between prediction and planning modules also introduces latency and computational inefficiency, making it challenging to meet the demands of real-time operation.
-Our key insight is that modeling the joint distribution of future trajectories for both the ego vehicle and surrounding agents can unify prediction and planning into a single, end-to-end framework.
-In this approach, the ego trajectory, informed by the desired goal location, represents the plan, while the trajectories of surrounding agents serve as predictions.
-Additional planning constraints can be imposed on the ego trajectory as needed.
-When trained on large-scale human driving datasets like the Waymo Open Motion Dataset (WOMD) [15], this joint distribution naturally captures the complex interactive behaviors observed in real-world scenarios.
-Diffusion models [51, 20] offer a promising method to sample from this high-dimensional, multimodal joint distribution of ego and surrounding agents’ trajectories.
-They have demonstrated success in generating high-quality images and videos [39, 21] or robot trajectories [38, 27].
-Moreover, their ability to model conditional distributions makes them well-suited for trajectory planning applications, as it allows for the incorporation of crucial context for planning and prediction, including trajectory history, map information, ego vehicle’s goal locations, etc.
-Compared to transformers, diffusion-based method also supports controllable generation to accommodate additional requirements solely at test time through guided sampling, without any additional training.
-However, diffusion models typically require many sampling steps to obtain high-quality samples, making it challenging to meet the demands of real-time operations on autonomous vehicles.
-To overcome this limitation, consistency models [52] are designed to generate data directly from noise in a single step or a small number of steps, significantly improving sampling efficiency without compromising quality.
-In this paper, we present an end-to-end predictive planner based on consistency models, designed to unify planning and prediction within a single data-driven framework.
-Our method samples from the conditional joint distribution of ego and surrounding agents’ future trajectories, given encoded map and trajectory history along with the ego vehicle’s goal configuration.
-While we use a transformer-based scene encoder [45] in our implementation, our approach is encoder-agnostic and can integrate with any suitable scene encoder.
-Trained on WOMD [15], our method ensures proactive trajectory planning by capturing complex interactions between agents observed in the real world.
-Additionally, during online testing, planning constraints are incorporated through guided sampling using our proposed alternating direction method for multi-objective guidance.
-Experiments demonstrate that our consistency model-based planner generates effective interactive behaviors with other agents, such as nudging and yielding, through a single inference, eliminating the need for iterative alternation between prediction and planning modules.
-It outperforms diffusion models by requiring fewer sampling steps while delivering superior trajectory quality.
-Compared to transformer-based models, it enables controllable generation through online guided sampling that significantly improves planning constraint satisfaction.
+Also, existing work attempts to incorporate the ego vehicle’s plan into the prediction module [18], but the initial ego plans may already be suboptimal if ignoring future interaction with others.
+In addition, the alternation between prediction and planning introduces computational inefficiency, making it challenging to meet the demands of real-time operation.
+Recently, there has been growing interest in multi-agent trajectory generation using purely data-driven approaches.
+Transformer-based methods have been widely adopted for multi-agent prediction tasks [35, 28].
+Diffusion models, with their strong ability to sample from high-dimensional multimodal distributions, have been adopted that improve the prediction performance [22].
+Compared to transformer-based models, diffusion models also support controllable generation through guidance methods, making them particularly suitable for scenario generation [19, 47, 8].
+However, diffusion models typically require many sampling steps to obtain high-quality samples, which becomes challenging to meet the demands of real-time operations on autonomous vehicles.
+To address this limitation, consistency models [40] are designed to generate data directly from noise in a single step or a small number of steps.
+This significantly improves sample efficiency and maintains high sample quality under limited computational budgets[12, 29].
+In this paper, we present a predictive planner for autonomous driving with a unified data-driven framework.
+Our key idea is to leverage a consistency model to sample the joint distribution of the ego and surrounding agents’ future trajectories, given the map, trajectory history, and the ego vehicle’s goal.
+Consistency models enable high-quality trajectory generation with only a few sampling steps, making them more suitable for online planning than standard diffusion models.
+Trained on WOMD [13], our method captures complex agent interactions and produces proactive trajectories observed in real-world driving.
+To further enforce multiple planning constraints, we introduce a novel online guided sampling approach inspired by ADMM [6].
+Our planner is also encoder-agnostic and can be integrated with any suitable scene encoder.
+Experiments show that our consistency model-based planner generates proactive behaviors, e.g., nudging and yielding, through a single forward pass without alternation between prediction and planning.
+It outperforms both diffusion models and an accelerated variant DDIM [37] with smoother, safer, and more efficient trajectories under limited computational budgets.
+Compared to transformer-based models, we enable controllable generation that significantly improves satisfaction of multiple trajectory constraints without affecting the original distribution.
 II Related Work
-For trajectory predictions, data-driven methods have shown promising results when trained on large-scale real-world human driving datasets such as Waymo Open Motion Dataset (WOMD) [15], INTERACTION Dataset [62], Argoverse [8] or nuScenes [7].
-Graph Neural Networks have been used to model the semantic information in traffic scenarios for interactive prediction [23].
-Generative models like Variational Autoencoders and Diffusion Models have also been utilized for multimodal trajectory prediction tasks [43, 27].
-Recently, transformers [56], particularly Motion Transformers (MTR) [45, 46, 34] with encoder-decoder architectures, have achieved state-of-the-art results in trajectory prediction [58].
-In trajectory planning, the goal is to determine a feasible and efficient path for the ego vehicle, based on observations of the environment and surrounding agents.
-This task is often formulated as a numerical optimization problem, typically decomposed path planning and speed planning to facilitate real-time operation in dynamic environments [60, 61, 2].
+Traditionally, trajectory planning is often formulated as a numerical optimization problem, typically decomposed into path planning and speed planning to facilitate real-time operation in dynamic environments [2].
+Beyond standalone planning, preliminary effort have integrated trajectory predictions into planner design for collision avoidance [27] and proactive motion planning [3, 33, 38], but either struggle to scale with the number of agents [20] or compromise optimality due to a restrictive exploration of the solution search space [38].
 Game-theoretic approaches have been explored to model agent interactions [31], with efficient computation [14] that considers imperfect traffic agents [43], safe actions [42], and uncertainty reduction in the estimates of agent behaviors [17].
-Beyond standalone planning, preliminary effort has been made to integrate trajectory predictions into planner design for collision avoidance [31] and proactive motion planning [4, 44, 49].
-While the aforementioned methods have advanced trajectory prediction and planning performance in dense environments, to our best knowledge, none have attempted to unify these two to achieve human-like interactive behavior with computational efficiency.
-Instead, our work addresses this gap by presenting a data-driven framework for end-to-end joint prediction and planning with a single inference.
-Diffusion models [47, 51, 20] are generative models capable of producing high-quality samples from complex distributions, initially achieving success in image and video generations [39, 21].
-The Denoising Diffusion Probabilistic Model (DDPM) [20] gradually corrupts the training data with Gaussian noise and learns to reverse this process through iterative denoising with deep neural networks.
-These models commonly use U-Net [40] or transformer-based architectures [56] as backbones.
-Conditional sampling in diffusion models is enabled by techniques such as classifier guidance [13] and classifier-free guidance [19] However, DDPM requires numerous sampling steps to preserve the sample quality, resulting in significant computational costs.
-To address this inefficiency, Denoising Diffusion Implicit Models (DDIM) [48] introduces a deterministic generation process that balances sampling quality and speed, while progressive distillation [42] reduces the number of sampling steps by distilling a pre-trained diffusion model into a more efficient one.
-Consistency model [52, 50] takes a different approach, directly generating data from noise.
-They can be trained through distillation of a pre-trained diffusion model or a standalone consistency training and outperform distillation methods in one- and few-step sampling [52].
-Diffusion and consistency models also have shown success in robotics and autonomous systems for their expressiveness and ability to provide samples from complex distributions.
-In robotics, conditional diffusion models have been applied to policy learning [26, 1, 11], with consistency models following similar applications [14, 37].
-Diffusion models have also been adopted for trajectory optimization across various robotics tasks [33, 32].
-In autonomous driving, diffusion models have shown promising results in trajectory prediction [27, 57] and controllable traffic simulation or scenario generation [63, 59, 9, 24], capturing complex interactive behaviors in traffic scenarios.
-However, our work differs fundamentally from these efforts: we focus on developing a predictive trajectory planner with consistency models, leveraging guided sampling to incorporate planning constraints.
-Unlike scenario generation tasks, where sampling efficiency is less critical, our consistency model-based predictive planner needs significantly shorter sampling steps to deliver high-quality interactive behaviors compared to diffusion models.
+Recently, data-driven methods have shown promising results in trajectory predictions when trained on large-scale real-world human driving datasets such as Waymo Open Motion Dataset (WOMD) [13], nuScenes [7], etc.
+Transformer [44] based methods, such as Scene Transformer [28] and Motion Transformers (MTR) [34], use an encoder-decoder architecture and achieved great performances on Waymo Motion Prediction Challenge [45].
+They allow for multi-agent trajectory generation [28, 35] with potential for goal-conditioned trajectory planning.
+However, trajectory planning usually requires more controllable generation to satisfy constraints provided online, which remains challenging for transformer-based approaches.
+Diffusion models are powerful generative models capable of producing high-quality samples from complex distributions, initially gaining success in computer vision [36, 41, 16].
+The Denoising Diffusion Probabilistic Model (DDPM) [16] corrupts training data with Gaussian noise and learns to reverse this process through iterative denoising.
+Controllable generation in diffusion models can be achieved through classifier guidance [11] during sampling.
+As a result, Diffusion models have been applied to robot policy learning [21, 1, 10] and trajectory generation [26, 25, 46].
+In autonomous driving, diffusion model facilitates multimodal trajectory predictions [22] and controllable traffic scenario generation [47, 8, 19], capturing complex interactive behaviors.
+However, diffusion models typically require a large number of sampling steps to maintain sample quality, resulting in substantial computational overhead.
+To address this inefficiency, Denoising Diffusion Implicit Models (DDIM) [37] introduces a deterministic sampling process that balances speed and quality, while progressive distillation [32] reduces the number of sampling steps by distilling a pre-trained diffusion model into a more efficient one.
+Consistency model [40, 39] takes a different approach by directly generating data from noise in a single or a few steps.
+They outperform distillation methods [40] in image generation and have shown early success in robot policy learning [12, 29].
+In this work, we focus on building a goal-conditioned predictive planner using a consistency model.
+Unlike offline scenario generation, where sampling efficiency is less critical, our method produces high-quality interactive behaviors with significantly fewer sampling steps than diffusion models, making it well-suited for high-frequency online planning.
+Compared to diffusion-based predictors, which model the full multimodal future distribution, our approach focuses on goal-conditioned generation tailored for planning.
+In contrast to transformer-based models, which lack controllable generation at inference time, our approach enables multiple planning constraints, e.g., goal-reaching and control limits, to be satisfied simultaneously through an ADMM-inspired online guided sampling.
 III Problem Formulation
-We aim to plan the trajectory of the ego vehicle with a planning goal, simultaneously predicting the interactive behaviors of others.
-We assume the trajectory histories of traffic agents and map information are given and known.
+Assuming an ego vehicle’s goal is given from a high-level planner and the trajectory histories of all traffic agents and map information are known, we aim to plan the trajectory of the ego vehicle while simultaneously predicting the interactive behaviors of others.
 At the current timestep $t_{0}\in\mathbb{R}$, the ego agent’s state history over the past $H_{1}\in\mathbb{N}$ timesteps including the current state is given as $S_{e}^{\text{hist}}=(s_{e}^{t_{0}},s_{e}^{t_{0}-1},...,s_{e}^{t_{0}-H_{1}})\in\mathbb{R}^{(H_{1}+1)\times d_{s}}$, where $d_{s}$ is the dimension of the states, e.g.
 position, velocity, etc.
-Let $\mathcal{O}$ denote the set of other agents of all types such as vehicles, pedestrians, and cyclists.
+Let $\mathcal{O}$ denote the set of other agents of all types, such as vehicles, pedestrians, and cyclists.
 Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$.
 The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c$ points on each polyline and $d_{m}$ attributes per point.
-We use fixed dimensions $N_{\mathcal{O}}$ and $N_{m}$ for agents and map polylines respectively, which are chosen large enough to accommodate all scenarios with padding and masks applied to invalid entries.
-We assume the ego agent’s goal position is given by some high-level planner and is notated $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
+We pad and mask the invalid entries.
+We assume the ego agent’s goal position is given as $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
 Let $\mathcal{O}_{s}\subset\mathcal{O}$ denote the subset of surrounding agents that may interact with the ego agents.
@@
 IV Preliminaries
-IV-A Motion-Transformer Encoder
+IV-A Motion Transformer Encoder
 To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [34, 35].
-This transformer-based architecture effectively models scene context through a local attention mechanism, enabling efficient interaction modeling between agents and the road map while maintaining memory efficiency.
-It also introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually.
-For more detailed descriptions of this encoder, we refer readers to the original MTR paper [45].
+This transformer-based encoder effectively models scene context and agents’ interactions, and introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually.
 The MTR++ encoder [35] would be suitable for multi-agent settings, but its implementation is not yet available.
-More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to integrate with our consistency model-based predictive planner.
-We represent the encoded feature of the trajectory history and map with MTR encoder as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
+More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to be integrated.
+We represent this encoded feature of the trajectory history and map as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
 IV-B Consistency Model
@@
 It consists of a forward diffusion process and a reverse process.
-Let $\mathcal{D}$ be our dataset of trajectories.
-We first normalize all trajectories in $\mathcal{D}$ using the empirical mean and standard deviation computed across the entire dataset.
-Let $\mathcal{X}$ be the space of such normalized trajectories, and $p_{\text{data}}$ be the data distribution over $\mathcal{X}$.
-In the forward process, we first draw an initial sample $x_{1}$ from $p_{\text{data}}$.
-We then apply an increasing noise schedule ${\sigma_{1},\sigma_{2},...,\sigma_{T}}$ with $\sigma_{1}\approx 0$ to progressively corrupt $x_{1}$ through $T$ steps.
-Specifically, at each step $i$, we draw a noise sample $\epsilon$ from $\mathcal{N}(0,I)$ and set the corrupted data $x_{i}$ as:
-$\quad x_{i}=x_{1}+\sigma_{i}\cdot\epsilon,\quad i=1,2,...,T$ (1)
-We choose $\sigma_{T}$ large enough so that, when $x_{1}$ is repeatedly sampled from $p_{\text{data}}$ and corrupted via additive Gaussian noise, the empirical distribution of the resulting $x_{T}$ approximates $\mathcal{N}(0,\sigma_{T}^{2}I)$.
+Let $p_{\text{data}}$ be the data distribution with zero mean and unit variance over space $\mathcal{X}$.
+In the forward process, we first draw an initial sample $x_{1}$ from $p_{\text{data}}$, then gradually corrupt it with an increasing noise schedule ${\sigma_{1},\sigma_{2},...,\sigma_{T}}$ through $T$ steps, with $\sigma_{1}\approx 0$.
+At each step $i$, the corrupted data is $x_{i}=x_{1}+\sigma_{i}\cdot\epsilon$ with $\epsilon$ sampled from $\mathcal{N}(0,I)$.
+$\sigma_{T}$ is chosen large enough such that the empirical distribution of the resulting $x_{T}$ approximates $\mathcal{N}(0,\sigma_{T}^{2}I)$.
 Let $\mathcal{Y}$ be the space of conditioning information.
 In the reverse process, we aim to learn a consistency function $f_{\theta}$ with the parameter $\theta$ that maps a noisy trajectory sample $x_{i}$, condition $y\in\mathcal{Y}$, and the noise level $\sigma_{i}$ directly to the corresponding clean sample $x_{1}$ for each step $i=1,..,T$.
-This is achieved by choosing a specific function form of $f_{\theta}$:
-$f_{\theta}(x,y,\sigma)=c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(x,y,\sigma)$ (2)
-where $\sigma$ represents the noise level, and the differentiable function $c_{\text{skip}}$ and $c_{\text{out}}$ are chosen such that $c_{\text{skip}}(\sigma_{1})=1,c_{\text{out}}(\sigma_{1})=0$ when $\sigma=\sigma_{1}$ [28].
-This ensures that the boundary condition of $f_{\theta}$ is met: $f_{\theta}(x_{1},y,\sigma_{1})=x_{1}$.
-$F_{\theta}$ is usually approximated by the neural network using a U-Net [40] or a transformer [36] based architecture.
-For the consistency model training, we aim to enforce the consistency of the output from $f_{\theta}$ for adjacent sampling steps $i$ and $i+1$ to be both $x_{1}$.
-Let $\mathcal{P}[1,T-1]$ be a distribution over the integer set ${1,2,...,T-1}$.
-The consistency training minimizes the following loss function $\mathcal{L}_{\text{consistency}}$:
+This is achieved by choosing a specific function form of $f_{\theta}\coloneqq c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(x,y,\sigma)$, where $c_{\text{skip}}$ and $c_{\text{out}}$ are differentiable function with $c_{\text{skip}}(\sigma_{1})=1,c_{\text{out}}(\sigma_{1})=0$ [23], such that the boundary condition is met: $f_{\theta}(x_{1},y,\sigma_{1})=x_{1}$.
+$F_{\theta}$ is usually approximated by neural networks.
+For the consistency model training, we aim to enforce the consistency of the output from $f_{\theta}$ for adjacent sampling steps $i$ and $i+1$ to be both $x_{1}$, through the following loss $\mathcal{L}_{\text{consistency}}$:
 $\mathcal{L}_{\text{consistency}}=$
 $\mathbb{E}_{i\sim\mathcal{P}[1,T-1],\epsilon\sim\mathcal{N}(0,I)}\bigl{[}d(f_{\theta}(x_{i},y,\sigma_{i})-f_{\theta}(x_{i+1},y,\sigma_{i+1}))\bigr{]}$ (1)
-where $\mathcal{P}$ can be the uniform distribution or lognormal distribution, and $d$ is chosen to be a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [50].
-During the data generation process, we first draw a sample $x_{T}$ from $\mathcal{N}(0,\sigma_{T}^{2}I)$.
+where $\mathcal{P}[1,T-1]$ is lognormal distribution over the integer set $\{1,2,...,T-1\}$, and $d$ is a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [39].
+For data generation, we draw a sample $x_{T}$ from $\mathcal{N}(0,\sigma_{T}^{2}I)$.
 Then with the trained consistency model $f_{\theta}$, we perform iterative sampling by first predicting the approximate clean data $x_{1}^{[i]}$ and then sample $x_{i-1}$ with $\epsilon$ drawn from $\mathcal{N}(0,I)$:
 $x_{1}^{[i]}=f_{\theta}(x_{i},y,\sigma_{i}),\quad x_{i-1}=x_{1}^{[i]}+\sigma_{i-1}\epsilon$ (2)
-for $i=T,T-1,...,1$, until we obtain the final clean sample $x_{1}^{[1]}$.
+for $i=T,T-1,...,1$, until we obtain the clean sample $x_{1}^{[1]}$.
 V Methodology
-With the aforementioned notations and stated problem, our proposed method is summarized in Fig.
+Our method is summarized in Fig.
 1.
 We first use an MTR encoder to encode the agent’s trajectory history and map information.
-Then we use a consistency model that takes the conditional input of the ego agent’s planning goal and the MTR-encoded features and generates trajectory plans and predictions for the ego and surrounding agents, respectively.
-Additional planning constraints for the ego agent are achieved through guided sampling of consistency models.
+With the ego agent’s planning goal and the MTR-encoded features, a consistency model generates trajectory plans and predictions for the ego and surrounding agents, respectively.
+Multiple planning constraints for the ego agent are achieved through ADMM-inspired online guided sampling.
 V-A Data pre-processing
-For training diffusion models or consistency models, we usually need to normalize the input data.
-If we jointly model ego and surrounding agents’ future trajectories $(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$, using an ego-centric coordinate system introduces large variances in the data, particularly for surrounding agents whose positions vary significantly across scenarios.
-Thus the training performance is largely degraded.
-Instead, we apply a coordinate transformation $\Gamma$ similar to MTR++ [46] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$.
-This transformation provides us with data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ that has substantially reduced variance.
-We then compute the empirical mean and standard deviation of the transformed trajectories across the dataset and standardize them to zero mean and unit variance.
-To preserve the relative spatial relationships between agents, we collect the reference states consisting of each agent’s position at timestep $t_{0}$ as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$.
+We apply a coordinate transformation $\Gamma$ similar to MTR++ [35] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$.
+The transformed data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ has substantially reduced variance.
+We then standardize the transformed trajectories to zero mean and unit variance.
+To preserve the relative spatial relationships among agents, the reference states are collected as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$ consisting of each agent’s position at timestep $t_{0}$.
 V-B Consistency Model Training
-Let $\mathcal{X}$ denote the space of transformed future trajectories of the ego vehicle $\Gamma(S_{e}^{\text{future}})$ and surrounding agents $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec.
+(a) A proactive nudging behavior from the ego car to switch a lane.
+(b) A safe yielding behavior for the ego car to make a turn.
+Figure 2: For each scenario, upper row: interactive behavior from our predictive planner; lower row: Ground truth driving behaviors.
+The ego agent’s trajectory is in green annotated by “E” and the surrounding agents’ trajectories are in orange annotated by“S”.
+The other agents’ trajectories (always shown with ground truth) are in blue.
+Let $\mathcal{X}$ denote the space of transformed future trajectories $\Gamma(S_{e}^{\text{future}})$ and $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec.
 V-A.
 Let $\mathcal{Y}$ denote the space of conditional inputs, containing encoded historical trajectories $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$, map features $\text{MTR}(M)$ from Sec.
-IV-A, ego agent’s goal states $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec.
+IV-A, ego agent’s planning goal $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec.
 V-A.
-Given a planning goal and environment context as the condition $y\in\mathcal{Y}$, we leverage the consistency model [52, 50] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$.
+Given a sample $y\in\mathcal{Y}$, we leverage the consistency model [40, 39] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$.
 Each trajectory sample represents one possible joint future behavior of the ego vehicle and surrounding agents.
-For training our predictive planner, we jointly train the MTR encoder and the consistency model together in one step with a hybrid loss function, which is constructed as a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq.
+For training our predictive planner, we jointly train the MTR encoder and the consistency model together with a hybrid loss function, which is a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq.
 (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec.
@@
 where $\omega_{0}$ and $\omega_{1}$ control the weight of these two losses.
-V-C Guided Sampling
-When generating trajectories with the trained consistency model as described in Sec.
-IV-B, to impose planning constraints on the ego vehicle’s future trajectory, we present a novel guided sampling approach similar to classifier guidance [13].
-Importantly, this guidance is applied solely during sampling at test time, without requiring any modification to the training procedure.
-Assuming there are $N_{c}\in\mathbb{N}$ planning constraint $c_{j}$ to minimize, $j=1,2,...,N_{c}$ and drawing inspiration from classifier guidance [13], we can perform a gradient descent step on the predicted $x_{1}^{[i]}$ during each sampling step $i$ in Eq.
-(4):
+V-C Planning Constraints Construction
+To make the sampled trajectories suitable and smooth for planning use, we apply several constraints on the ego trajectories. Assuming the vehicle has the following dynamics
+$\dot{p}_{x}=v\cos\theta,\quad\dot{p}_{y}=v\sin\theta,\quad\dot{v}=a,\quad\dot{\theta}=\omega$ (4)
+where $p_{x}$ and $p_{y}$ are positions, $v$ is the linear speed and $\theta$ is the orientation.
+The control input is the acceleration $a$ and angular speed $\omega$.
+We only use $p_{x}$ and $p_{y}$ from the future states $S_{e}^{\text{future}}$ to construct our planning constraint function $c_{i}$, since other states like $v$ or $\theta$ in the dataset may not satisfy the dynamics equations in Eq.
+(4) and are also quite noisy.
+Using the differential flatness for dynamics in Eq.
+(4), we infer the control input $a$ and $\omega$ as $a=\frac{\dot{p}_{x}\ddot{p}_{x}+\dot{p}_{y}\ddot{p}_{y}}{\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}},\omega=\frac{\dot{p}_{x}\ddot{p}_{y}-\dot{p}_{y}\ddot{p}_{x}}{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}$ through finite differencing with inevitable noises.
+With $p_{x},p_{y}$ and the derived $a$ and $\omega$, we consider three types of planning constraints $c$ to minimize as follows.
+$c_{\text{goal}}=\sqrt{(p_{x}^{t_{0}+H_{2}}-p_{x,goal})^{2}+(p_{y}^{t_{0}+H_{2}}-p_{y,goal})^{2}}$
+$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
+$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (5)
+$c_{\text{goal}}$ enforces goal-reaching for the ego vehicle at the final timestep $t_{0}+H_{2}$.
+$c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints over the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$.
+These constraints are inherently in conflict.
+For example, only moving the trajectory’s final state closer to the goal can increase acceleration violations of the car
+V-D Guided Sampling
+To simultaneously enforce multiple planning constraints, like in Eq.
+(V-C), on the trajectory samples without affecting the original distribution, we present a novel ADMM-inspired guided sampling approach, which is applied solely during sampling at test time, without requiring any modification to the training procedure.
+Assuming there are $N_{c}\in\mathbb{N}$ planning constraint $c_{j}$ to minimize, $j=1,2,...,N_{c}$, standard classifier-guidance [11] performs a vanilla gradient descent step on the predicted $x_{1}^{[i]}$ during each sampling step $i$ in Eq.
+(2) as follows:
 $x_{1}^{[i,1]}=x_{1}^{[i]}-\sum^{N_{c}}_{j=1}\alpha_{j}\nabla c_{j}$ (6)
@@
 (2).
-However, optimizing multiple constraints simultaneously presents a significant challenge in finding a suitable stepsize $\alpha$ efficiently, particularly when constraints may conflict with each other.
-To address this challenge, inspired by the Alternating Direction Method of Multipliers (ADMM) [6], we propose a novel alternating direction method that only optimizes one constraint at a time during each sampling step $i$:
-$x_{1}^{[i,1]}=x_{1}^{[i]}-\alpha_{1}\nabla c_{1}$
-$x_{1}^{[i,2]}=x_{1}^{[i,1]}-\alpha_{2}\nabla c_{2}$
-$x_{1}^{[i,N_{c}]}=x_{1}^{[i,N_{c}-1]}-\alpha_{N_{c}}\nabla c_{N_{c}}$ (7)
-In our approach, we sequentially optimize each constraint $c_{j}$ with corresponding step sizes $\alpha_{j}$ during each gradient descent iteration.
-While our selection of the optimization order and relatively small step sizes is based on empirical observations, this strategy has demonstrated effective convergence in practice, which is difficult to achieve with the standard guidance method with gradient descent in Eq.
+However, optimizing multiple constraints simultaneously with conflicts presents a challenge in finding a suitable stepsize $\alpha$.
+To address this, we propose a novel alternating direction method inspired by ADMM [6] that sequentially applies gradient step for each constraint during the sampling step $i$:
+$x_{1}^{[i,j+1]}=x_{1}^{[i,j]}-\alpha_{j}\nabla c_{j},\quad\text{for}\quad j=1,...,N_{c}$ (7)
+We find that this alternating update strategy demonstrates effective convergence and empirically leads to lower violations of multiple conflicting constraints, compared to standard guided sampling with vanilla gradient descent in Eq.
 (6).
-The convergence of alternating direction methods, such as ADMM, is generally guaranteed under conditions including closed, proper, and convex functions and appropriately chosen step sizes [6].
-V-D Planning Constraints Construction
-We assume that the ego vehicle’s dynamics are as follows:
-$\dot{p}_{x}=v\cos\theta,\quad\dot{p}_{y}=v\sin\theta,\quad\dot{v}=a,\quad\dot{\theta}=\omega$ (8)
-where $p_{x}$ and $p_{y}$ are positions, $v$ is the linear speed and $\theta$ is the orientation. The control input is the acceleration $a$ and angular speed $\omega$.
-Suppose we only use $p_{x}$ and $p_{y}$ from ego agent’s future states $S_{e}^{\text{future}}$ to construct our planning constraint function $c_{i}$, since other states like $v$ or $\theta$ may not satisfy the dynamics equations in Eq.
-(8) and are also quite noisy.
-Using the property of differential flatness for dynamics in Eq.
-(8), we can infer $v$ and $\theta$ with $p_{x}$ and $p_{y}$ as follows:
-$v=\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}},\quad\theta=\arctan(\frac{\dot{p}_{y}}{\dot{p}_{x}})$ (9)
-Then we can further infer the control input $a$ and $\omega$:
-$a=\frac{\dot{p}_{x}\ddot{p}_{x}+\dot{p}_{y}\ddot{p}_{y}}{\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}},\quad\omega=\frac{\dot{p}_{x}\ddot{p}_{y}-\dot{p}_{y}\ddot{p}_{x}}{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}$ (10)
-Note that, $a$ and $v$ are computed through finite differencing of $p_{x}$ and $p_{y}$ which will inevitably incur some noises.
-With $p_{x},p_{y}$ and the derived $a$ and $\omega$, we consider three types of planning constraint $c$ to minimize as follows.
-$c_{\text{goal}}=\sqrt{(p_{x}^{t_{0}+H_{2}}-p_{x,goal})^{2}+(p_{y}^{t_{0}+H_{2}}-p_{y,goal})^{2}}$
-$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
-$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (11)
-$c_{\text{goal}}$ represents a goal-reaching constraint such that the ego vehicle needs to reach a pre-specified goal $(p_{x,\text{goal}},p_{y,\text{goal}})$ at the final timestep $t_{0}+H_{2}$.
-$c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints that compute the average control limit violation along the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$.
+It also simplifies step size tuning.
 VI Experiments
 VI-A Experiment setup
+Figure 3: Planning with different goal locations using consistency models and guided sampling.
+TABLE I: Planning Trajectory Realism
+Method minADE ($m$) minFDE ($m$) Step
+Transformer 1.237 2.796 1
+Diffusion-10 0.488 0.532 10
+Diffusion-4 0.545 0.711 4
+DDIM 2.749 2.304 4
+Ours 0.497 0.173 4
+Ours (guided) 0.490 0.008 4
+TABLE II: Planning Trajectory Quality and Constraint Violations for Different Methods
+Method Planning Trajectory Quality Planning Constraint Violations
+Angle changes Path length Curvature Collision rate Goal reaching Acceleration limit $\omega$ limit
+[…]
+Transformer 0.465 55.885 0.599 15.0 8.111 0.438 1.119
+Diffusion-10 0.639 67.899 1.392 1.7 2.410 35.609 2.248
+Diffusion-4 0.716 61.407 1.379 2.0 3.171 19.812 2.566
+DDIM 0.673 135.55 1.755 18.6 14.440 175.602 4.545
+Ours 0.670 55.373 1.107 0.8 0.361 4.806 2.064
+Ours (guided) 0.487 55.137 0.556 0.8 0.026 0.281 0.519
+TABLE III: Increased constraint violation (%) for vanilla gradient descent compared to the ADMM-inspired method.
+Gradient Step Goal reaching Acceleration limit $\omega$ limit
+50 10% 26% 8%
+100 9% 35% 8%
+150 12% 41% 9%
 VI-A1 Dataset
-We evaluate our method on the Waymo Open Motion Dataset (WOMD) [15], which contains 486k traffic scenarios in the training set and 44k scenarios in the validation set that we can test on.
-Each scenario provides road graph information and trajectories of diverse agents including vehicles, pedestrians, and cyclists, with up to 8 agents specifically marked for trajectory prediction.
-For each agent, this dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals.
-This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps, respectively.
-For each scenario, with the trajectory information from all agents including vehicles, pedestrians, and cyclists as $S_{\mathcal{O}}$, we only select vehicle agents as the ego agent with trajectories denoted by $S_{e}$.
-We then identify the surrounding agents, with trajectories denoted as $S_{\mathcal{O}_{s}}$, by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold.
+We evaluate our method on the Waymo Open Motion Dataset (WOMD) [13], which contains 486k traffic scenarios in the training set and 42k scenarios in the interactive validation set to test on.
+For each agent, including vehicles, pedestrians, and cyclists, the dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals.
+This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps.
+We identify the surrounding agents with trajectories $S_{\mathcal{O}_{s}}$ by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold.
 This enables us to focus our predictive planner on the most relevant interactive behaviors between the ego vehicle and its immediate surroundings.
@@
 VI-A2 Model Details
-The MTR encoder consists of 6 transformer layers with a dimension of 256 to encode agent trajectory histories and map information into $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M)$.
-For detailed encoder architecture specifications, we refer readers to [45].
-The conditional input $y$ for the consistency model is constructed by further encoding three components: the obtained features $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M)$, the ego agent’s goal $s_{e}^{g}$, and the reference states $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$, using 3-layer MLPs with 256 neurons per layer.
-To model the consistency model function $F_{\theta}$, we adopt the U-Net [40] with the model dimension of 128 as the backbone.
-For the consistency model, we choose $T=5$ which indicates 4 sampling steps in the forward and reverse process.
-Follow the original paper [50], we set the noise schedule with $\rho=6$, $\sigma_{1}=0.002,\sigma_{T}=80$, and then $\sigma_{i}=\left(\sigma_{0}^{1/\rho}+\frac{i-1}{T-1}(\sigma_{T}^{1/\rho}-\sigma_{0}^{1/\rho})\right)^{\rho},i=1,...,T$.
+The MTR encoder consists of 6 transformer layers with a dimension of 256 to encode agent trajectory histories and map information [34].
+The consistency model’s conditional embedding uses 3-layer MLPs with 256 neurons per layer.
+The consistency function $F_{\theta}$ is modeled with a U-Net [30] with a dimension of 128.
+We choose $T=5$, i.e., 4 sampling steps in the forward and backward process for the consistency model.
+We set the noise schedule with $\rho=6$, $\sigma_{1}=0.002,\sigma_{T}=80$, and $\sigma_{i}=\left(\sigma_{0}^{1/\rho}+\frac{i-1}{T-1}(\sigma_{T}^{1/\rho}-\sigma_{0}^{1/\rho})\right)^{\rho},i=1,...,T$ as  [39].
 During guided sampling, we perform 100 gradient steps of Eq.
-(V-C) at each sampling step, and set the coefficient for $c_{\text{goal}},c_{\text{acc}}$ and $c_{\omega}$ to be $2\times 10^{-5},3\times 10^{-6},5\times 10^{-7}$, respectively.
-Figure 2: Ego vehicle’s planned trajectory (Green) in a left-turn scenario compared to the Groundtruth (Red).
-Figure 3: Planning with different goal locations using consistency models and guided sampling.
-TABLE I: Planning Trajectory Compared to the Human Groundtruth
-Method minADE $\downarrow$ minFDE $\downarrow$
-Transformer 0.944 3.239
-DDPM-10 0.346 0.377
-DDPM-4 0.363 0.498
-DDIM-4 2.126 2.559
-Consistency 0.309 0.218
-Consistency (guided) 0.303 0.016
-(a) A proactive nudging behavior from the ego car to switch a lane.
-(b) A yielding behavior for the ego car to make a turn.
-Figure 4: For each scenario, upper row: interactive behavior from our predictive planner; lower row: Ground truth driving behaviors.
-The ego agent’s trajectory is in green annotated by “E” and the surrounding agents’ trajectories are in orange annotated by“S”.
-The other agents’ trajectories (always shown with ground truth) are in blue.
+(7) at each sampling step, and set the coefficient for $c_{\text{goal}},c_{\text{acc}}$ and $c_{\omega}$ to be $2\times 10^{-5},3\times 10^{-6},5\times 10^{-7}$, respectively.
 VI-A3 Methods and Baselines
 We compare our consistency model-based approach against several baselines.
-Currently, we only provide the evaluation of the open-loop trajectory generation performance of those methods, focusing on the behavior through a single inference step.
+Currently, we provide the evaluation of the open-loop trajectory generation performance of those methods, focusing on interactive behavior through a single inference.
 Transformer
-A model based on MTR’s encoder-decoder architecture [45] to generate trajectory only for ego vehicle.
+This transformer baseline uses MTR [34] to generate trajectory only for ego vehicle.
 It has a larger model size for the decoder (512 vs.
-our 128), but doesn’t incorporate ego vehicle’s goal $s_{e}^{g}$ as input.
-DDPM
-This diffusion baseline uses the same MTR encoder and MLP architecture as ours for condition encoding but adopts DDPM [20] for trajectory sampling of ego and surrounding agents.
-Note that it differs from consistency models in both its forward diffusion process and training loss function.
-We evaluate two variants: one trained and sampled with 4 steps (DDPM-4), and another trained and sampled with 10 steps (DDPM-10).
+our 128), but doesn’t incorporate the ego vehicle’s goal $s_{e}^{g}$ as input.
+Diffusion-4 and Diffusion-10
+These diffusion baselines use the same MTR encoder and goal embedding as our approach but adopt DDPM [16] for trajectory sampling.
+We evaluate two variants: one trained and sampled with 4 steps (Diffusion-4), and another with 10 steps (Diffusion-10).
 DDIM
-This baseline extends DDPM-10 by incorporating the more efficient DDIM sampling strategy [48].
-While using the same training process as DDPM-10, it employs 4-step DDIM sampling during inference, denoted as DDIM-4.
-Consistency
-Our proposed consistency model is trained and sampled with 4 steps.
-Consistency (guided)
-Our proposed consistency model is trained and sampled with 4 steps, with guided sampling.
+This baseline aims to accelerate the pretrained Diffusion-10 model with only 4 sampling steps in trajectory generation, with an efficient DDIM sampling strategy [37].
+Ours and Ours guided
+These are our proposed consistency model trained and sampled with 4 steps, with and without the proposed guided sampling.
 VI-A4 Training
-We conduct distributed training with a total batch size of 100 across 4 NVIDIA L40 GPUs, each with a batch size of 25.
+We conduct distributed training with a total batch size of 100 across 4 NVIDIA L40 GPUs.
 For consistency and diffusion models, we use the Adam optimizer [24] with a learning rate of $8\times 10^{-5}$ and train for 50 epochs.
@@
 VI-B Experiment results and analysis
-In this section, we evaluate our consistency model-based planner from several aspects.
-First, we assess its planning performance and trajectory quality compared to several baseline methods.
-Then we take a closer look at interactive scenarios and present examples of proactive nudging and yielding behaviors of our predictive planner.
-Furthermore, we show how online guided sampling improves constraint satisfaction in planning - the controllability not provided with transformer-based approaches.
-Finally, we discuss the computational time during inference for our presented methods.
-VI-B1 Trajectory Planning Performance
+VI-B1 Interactive Behavior
+We demonstrate that our predictive planner can generate safe and effective interactions with other road users with a single inference of the consistency model, without alternation between prediction and planning.
 In Fig.
-2, we compare planned trajectories generated by different methods (shown in green) against the ground truth from the dataset (shown in red).
-The scenario requires the ego vehicle to make a left turn and then drive straight to reach its goal.
-According to Fig.
-2, DDPM-4 reaches the goal but produces a noisy trajectory, likely due to insufficient sampling steps for diffusion models to generate high-quality trajectories.
-While DDPM-10’s increased sampling steps yields slightly smoother trajectories, it comes at the cost of longer computation time.
-DDIM-4 attempts to accelerate DDPM-10 using only 4 sampling steps, but fails to generate goal-reaching trajectories.
-The transformer with a larger model size generates a trajectory close to the groundtruth but misses the exact goal location.
-In contrast, our consistency model generates smooth trajectories that both align with the ground truth and precisely reach the target location.
-In Table I, we quantitatively evaluate the planning performance for the ego vehicle compared to the human groundtruth through metrics from Waymo’s Motion Prediction Challenge [58]: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) on the entire test dataset.
-Here we view these metrics as an evaluation of how closely our generated trajectories align with human driving behavior under the same environmental context and goal conditions.
-The results align with our qualitative observations: our consistency model achieves the lowest minADE and minFDE, indicating that it captures human driving patterns well.
+2, we show a proactive nudging and a safe yielding behavior of our planner, with the ego agent (“E”) trajectory in green, the surrounding agents (“S”) in orange, and other agents in blue.
+In Fig.
+2a, our planner directs the ego vehicle to accelerate and nudge into the left lane, predicting that the vehicle in the left lane will decelerate to accommodate the lane change.
+While in the ground truth, the rightmost vehicle eventually turns right instead of going straight, this discrepancy is expected for an 8-second future prediction.
+In Fig.
+2b, the planner instructs the ego vehicle to wait until the pedestrian has fully crossed before making the right turn.
+VI-B2 Planning Performance under Limited Computational Budget
+Figure 4: Constraint violation losses with respect to gradient steps during the final sampling step of the consistency model.
+Vanilla gradient descent struggles with three conflicting constraints, while the ADMM-inspired method optimizes smoothly.
+In Table I, we quantitatively evaluate the realism of the planned trajectory for the ego vehicle through metrics with human groundtruth: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) [45] over 6 sampled trajectories.
+The results show that our consistency model, with only 4 sampling steps, captures realistic human driving patterns comparable to Diffusion-10.
 Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints.
-In Table II, we evaluate the trajectory quality using three metrics that characterize driving behavior: 1) angle changes, which measure the smoothness of directional transitions, 2) path length, which quantifies the efficiency of the chosen route, and 3) curvature, which assesses the sharpness of turns.
-Lower values in these metrics generally indicate smoother and more efficient trajectories that align with natural driving patterns.
-Our consistency model achieves superior trajectory quality compared to all diffusion-based methods, including DDPM-10 despite its additional sampling steps.
-while the transformer method has slightly lower angle changes and curvature, this is expected given its substantially larger model size (512 vs.
+In Table II, we evaluate the trajectory quality using four metrics 1) angle change as the smoothness of directional transitions, 2) path length as the efficiency of the chosen route, 3) curvature as the sharpness of turns, and 4) collision rate ($2m$ threshold) as safety.
+Lower values in these metrics indicate smoother, safer, and more efficient trajectories that align with natural driving patterns.
+Our consistency model achieves superior trajectory quality compared to all Diffusion-based methods, including Diffusion-10, despite its additional sampling steps.
+While the transformer method has slightly lower angle changes, this is expected given its substantially larger model size (512 vs.
 our 128 dimensions).
-TABLE II: Planning Trajectory Quality
-Method Angle changes $\downarrow$ Path len $\downarrow$ Curvature $\downarrow$
-Transformer 0.419 59.621 0.560
-DDPM-10 0.660 87.801 3.444
-DDPM-4 0.729 79.731 2.211
-DDIM-4 0.683 157.78 2.012
-Consistency 0.636 58.060 1.174
-Consistency (guided) 0.445 57.564 0.610
 Furthermore, we demonstrate the flexibility of our approach in choosing goal location in Fig.
 3.
-Our consistency model maintains high-quality trajectory planning when targeting novel goal locations, even when they are not present in the dataset.
-The current transformer method with MTR doesn’t support the selection of goal locations.
-VI-B2 Interactive Behavior
-A key advantage of our predictive planner is its ability to ensure safe and effective interactions with other road users in traffic.
-Our planner demonstrates the capability for generating both proactive behaviors as shown in Fig.
-4, such as nudging for lane switching and yielding to pedestrians.
-The ego agent (labeled “E”) trajectory is in green and the surrounding agents (labeled “S”) trajectories are in orange.
-Other agent’s trajectories are shown in blue.
-These learned behaviors are similar to ground truth human drivers and ensure efficient navigation and safe interaction between the ego vehicle and other road users.
-Crucially, all planning and prediction are achieved through a single inference of the consistency model.
+Our consistency model maintains high-quality trajectory planning when targeting novel goal locations not present in the dataset.
+The current transformer method (MTR) doesn’t support goal selections.
+With the same encoder, the computational time for diffusion-based methods directly depends on the number of sampling steps, shown in Table I.
+Diffusion-4 and DDIM have the same sampling step as our model, but fail to match our performance in trajectory generation across all metrics, as evidenced in Table I and Table II.
+The guided sampling, introducing a little computational overhead ($\approx 15\%$), provides higher quality trajectory generation and enables controllable planning.
+Even Diffusion-10, despite its more sampling steps, cannot achieve a comparable performance of trajectory generation to our approach in Table II.
+VI-B3 Ablation Study on ADMM-inspired Guided Sampling
+We compare our ADMM-inspired method (ADMM for short) in Eq.
+(7) with the vanilla gradient descent (GD for short) in Eq.
+(6), which simultaneously optimizes all planning constraints.
+To illustrate the advantage of our method in hyperparameter tuning, we sweep through all combinations of step sizes ($\alpha=1e^{-4},1e^{-5},1e^{-6}$) across three constraints (27 combinations total), using 50, 100, and 150 gradient steps over 200 scenarios.
+TABLE III shows that GD results in significantly higher violations than ADMM on average; e.g., at 100 steps, GD has 9% higher goal-reaching, 35% higher acceleration, and 8% higher angular speed violations.
+The lowest sums of violations achieved are 1.45, 1.16, and 0.71 for 50, 100, and 150 steps, respectively.
+Increasing gradient steps reduces constraint violations but with computational tradeoffs.
 In Fig.
-4a, we show a proactive nudging scenario.
-Here, the ego vehicle needs to merge into the left lane while a nearby vehicle occupies that lane.
-Our planner directs the ego vehicle to accelerate and nudge into the left lane, predicting that the vehicle in the left lane will decelerate to accommodate the lane change.
-This nudging behavior demonstrates the proactive planning capability of our end-to-end, data-driven predictive planner, which is challenging to achieve using modular methods.
-While in the ground truth, the left car doesn’t decelerate that much and the rightmost vehicle eventually turns right, this discrepancy is expected given that we currently generate 8-second future trajectories through a single inference.
+4, we visualize how the violation loss evolves with ADMM (blue) and GD (red) methods in the last guided sampling step, measured across 200 scenarios and with the same step size finally chosen.
+As discussed in Sec.
+V-D, these three constraints are in conflict with each other.
 In Fig.
-4b, a yielding behavior is presented.
-The ego vehicle intends to make a right turn but predicts that a pedestrian will continue crossing the road.
-As a result, the planner instructs the ego vehicle to wait until the pedestrian has fully crossed, which is highly aligned with the ground truth.
-This yielding behavior ensures safe interaction between the ego vehicle and other road users, highlighting the planner’s safe operation in dynamic environments.
-VI-B3 Guided Sampling
-The guided sampling approach described in Sec.
-V-C and Sec.
-V-D allows us to enforce planning constraints during inference time, providing a key advantage over transformer-based methods which lack this capability.
-We quantitatively evaluate all methods against three planning constraints from Eq.
-(V-D): goal reaching, acceleration limits, and orientation speed limits.
-As shown in Table III, our base consistency model already achieves more accurate goal-reaching and smoother behaviors than other diffusion-based approaches.
-With guided sampling, these planning constraint violations of the consistency model further decrease by a factor of 4, achieving the lowest values among all methods.
-Note that this improvement requires no model retraining and only applies during inference.
-TABLE III: Planning Constraint Violations
-Method Goal Reaching $\downarrow$ Acc limit $\downarrow$ $\omega$ limit $\downarrow$
-Transformer 9.256 0.570 1.029
-DDPM-10 12.669 80.581 3.307
-DDPM-4 13.873 56.170 3.323
-DDIM-4 22.933 218.823 5.270
-Consistency 0.618 5.754 2.050
-Consistency (guided) 0.127 0.462 0.599
-VI-B4 Computational time discussions
-All methods in our evaluation support batch sampling.
-As shown in Table IV, we evaluate their inference times using a batch of 30 scenarios with 125 ego agents on an NVIDIA GeForce RTX 3060.
-While the transformer achieves the fastest inference time, it lacks our method’s key capabilities: online controllable generation and sampling from multimodal joint distribution.
-Our consistency model achieves a similar inference time with DDPM-4 and DDIM-4 due to the same model dimensionality, and sampling steps.
-However, DDPM-4 and DDIM-4 fail to match our consistency model’s performance in trajectory generation across all metrics, as evidenced in Fig.
-2, Table I, Table II, and Table III.
-The guided sampling, introducing a little computational overhead, provides higher quality trajectory generation and enables controllable planning.
-Even DDPM-10, despite its increased sampling steps and higher computational time, cannot achieve a comparable performance of trajectory generation to our approach.
-TABLE IV: Computational Time Comparison for Inference
-Method Inference Time (s)
-Transformer 1.910
-DDPM-10 3.352
-DDPM-4 2.226
-DDIM-4 2.118
-Consistency 2.301
-Consistency (Guided) 2.635
-VII Limitations
-Although our proposed method significantly improves the efficiency of generating human-like trajectories through consistency models and joint prediction and planning, it may not yet meet the demands of real-time operation, especially in highly dynamic environments that require high-frequency planning.
-Future work could leverage model pruning or distillation techniques to further improve the inference efficiency.
-Besides, while in the current paper, we test open-loop trajectory generation to focus on evaluating the model performance of a single inference step, in the future we will extend our method to the closed-loop case with a receding horizon fashion, which is expected to further enhance the interactive planning.
-VIII Conclusions
-In this paper, we present a novel data-driven approach that unifies trajectory prediction and planning using consistency models.
-By modeling the joint distribution of ego and surrounding agents’ trajectories, our method enables efficient and interactive planning, such as proactive nudging and yielding, without the need for iterative alternation between prediction and planning modules.
-Trained on the Waymo Open Motion Dataset, the proposed consistency model generates human-like trajectories while its guided sampling capability naturally incorporates additional planning constraints during online trajectory generations.
-Experimental results demonstrate that our approach achieves superior trajectory quality and interactive behavior compared to existing methods while requiring fewer sampling steps.
-This framework also has the potential to be extended to broader robot decision-making tasks in interactive scenarios.
-Looking forward, we aim to optimize this method for real-time operation and extend it to closed-loop operation with receding horizon planning.
-Additionally, we plan to incorporate LLMs as high-level planners to interpret diverse contextual information and enhance decision-making in challenging scenarios.
+4, GD experiences large oscillations in acceleration violations (middle plot), highlighting its struggles with optimizing conflicting objectives.
+In contrast, our ADMM method smoothly and effectively reduces all three violations.
+VII Conclusion
+We propose a consistency model-based predictive planner that generates safe and highly interactive behavior, such as nudging and yielding, for autonomous vehicles.
+Compared to diffusion models and their accelerated variants like DDIM, our method generates smoother, safer, and more efficient paths under a limited computational budget.
+To satisfy multiple planning constraints online, we introduce a novel ADMM-inspired approach used in guided sampling for consistency models.
+Future work will explore the integration of LLMs as high-level planners for closed-loop decision-making in long-tailed challenging scenarios.

### v3

@@
 Evaluated on the Waymo Open Motion Dataset (WOMD), our method enables proactive behavior such as nudging and yielding, and also demonstrates smoother, safer, and more efficient trajectories and satisfaction of multiple constraints under a limited computational budget.
+The project website is at https://anjianli21.github.io/projects/predictive_planner/.
 I Introduction
@@
 This typically requires a prediction module to anticipate other agents’ future trajectories and a planning module to generate trajectories for the ego vehicle.
-Traditionally, these two modules usually operate in a decoupled, alternating manner [27, 9], which introduces fundamental limitations.
+Traditionally, these two modules usually operate in a decoupled, alternating manner [9, 29, 46, 48], which introduces fundamental limitations.
 Some approaches integrate data-driven predictions into an optimization-based planner that is inherently reactive rather than proactive [16].
@@
 To address this limitation, consistency models [44] are designed to generate data directly from noise in a single step or a small number of steps.
-This significantly improves sample efficiency and maintains high sample quality under limited computational budgets[12, 29].
+This significantly improves sample efficiency and maintains high quality under limited computational budgets[12, 32].
 In this paper, we present a predictive planner for autonomous driving with a unified data-driven framework.
@@
 Our planner is also encoder-agnostic and can be integrated with any suitable scene encoder.
-Experiments show that our consistency model-based planner generates proactive behaviors, e.g., nudging and yielding, through a single forward pass without alternation between prediction and planning.
+Experiments show that our consistency model-based planner generates proactive behaviors, e.g., nudging and yielding, without alternation between prediction and planning.
 It outperforms both diffusion models and an accelerated variant DDIM [41] with smoother, safer, and more efficient trajectories under limited computational budgets.
@@
 II Related Work
-Traditionally, trajectory planning is often formulated as a numerical optimization problem, typically decomposed into path planning and speed planning to facilitate real-time operation in dynamic environments [2].
-Beyond standalone planning, preliminary effort have integrated trajectory predictions into planner design for collision avoidance [27] and proactive motion planning [3, 33, 38], but either struggle to scale with the number of agents [20] or compromise optimality due to a restrictive exploration of the solution search space [38].
+Traditionally, trajectory planning is often formulated as a numerical optimization problem, typically decomposed into path planning and speed planning to facilitate real-time operation in dynamic environments [2, 49].
+Beyond standalone planning, preliminary effort have integrated trajectory predictions into planner design for collision avoidance [29, 47, 22, 30] and proactive motion planning [3, 15, 34, 37], but either struggle to scale with the number of agents [21] or compromise optimality due to a restrictive exploration of the solution search space [42].
 Game-theoretic approaches have been explored to model agent interactions [35], with efficient computation [14] that considers imperfect traffic agents [51], safe actions [50], and uncertainty reduction in the estimates of agent behaviors [18].
@@
 Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$.
-The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c$ points on each polyline and $d_{m}$ attributes per point.
+The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c_{p}\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c_{p}$ points on each polyline and $d_{m}$ attributes per point.
 We pad and mask the invalid entries.
-We assume the ego agent’s goal position is given as $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
-Let $\mathcal{O}_{s}\subset\mathcal{O}$ denote the subset of surrounding agents that may interact with the ego agents.
-Given the information $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M,s_{e}^{g})$, for the future $H_{2}\in\mathbb{N}$ timesteps, we aim to plan the ego agent’s trajectory $S_{e}^{\text{future}}=(s_{e}^{t_{0}+1},s_{e}^{t_{0}+2},...,s_{e}^{t_{0}+H_{2}})\in\mathbb{R}^{H_{2}\times d_{s}}$, and predict the trajectories of surrounding agents $S_{\mathcal{O}_{s}}^{\text{future}}=(s_{o}^{t_{0}+1},s_{o}^{t_{0}+2},...,s_{o}^{t_{0}+H_{2}})_{o\in\mathcal{O}_{s}}\in\mathbb{R}^{N_{\mathcal{O}_{s}}\times(H_{2})\times d_{s}}$, where $N_{\mathcal{O}_{s}}=|\mathcal{O}_{s}|$.
+We assume the ego agent’s goal state is given as $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
+Let $\mathcal{O}_{s}\subset\mathcal{O}$ denote the subset of surrounding agents that may interact with the ego agent.
+Given the information $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M,s_{e}^{g})$, for the future $H_{2}\in\mathbb{N}$ timesteps, we aim to plan the ego agent’s trajectory $S_{e}^{\text{future}}=(s_{e}^{t_{0}+1},s_{e}^{t_{0}+2},...,s_{e}^{t_{0}+H_{2}})\in\mathbb{R}^{H_{2}\times d_{s}}$, and predict the trajectories of surrounding agents $S_{\mathcal{O}_{s}}^{\text{future}}=(s_{o}^{t_{0}+1},s_{o}^{t_{0}+2},...,s_{o}^{t_{0}+H_{2}})_{o\in\mathcal{O}_{s}}\in\mathbb{R}^{N_{\mathcal{O}_{s}}\times H_{2}\times d_{s}}$, where $N_{\mathcal{O}_{s}}=|\mathcal{O}_{s}|$.
 This joint prediction and planning naturally captures interactive behaviors between agents.
@@
 To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [38, 39].
-This transformer-based encoder effectively models scene context and agents’ interactions, and introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually.
+This transformer-based encoder effectively models scene context and agents’ interactions, and introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ [38] to train this encoder individually.
 The MTR++ encoder [39] would be suitable for multi-agent settings, but its implementation is not yet available.
@@
 $\mathbb{E}_{i\sim\mathcal{P}[1,T-1],\epsilon\sim\mathcal{N}(0,I)}\bigl{[}d(f_{\theta}(x_{i},y,\sigma_{i})-f_{\theta}(x_{i+1},y,\sigma_{i+1}))\bigr{]}$ (1)
-where $\mathcal{P}[1,T-1]$ is lognormal distribution over the integer set $\{1,2,...,T-1\}$, and $d$ is a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [39].
+where $\mathcal{P}[1,T-1]$ is lognormal distribution over the integer set $\{1,2,...,T-1\}$, and $d$ is a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+\delta^{2}}-\delta$ [43] where $\delta$ is a small constant.
 For data generation, we draw a sample $x_{T}$ from $\mathcal{N}(0,\sigma_{T}^{2}I)$.
@@
 V-A Data pre-processing
-We apply a coordinate transformation $\Gamma$ similar to MTR++ [35] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$.
+Instead of using an ego-centric coordinate frame, we apply a coordinate transformation $\Gamma$ similar to MTR++  [39] that maps each agent’s trajectory into its own local coordinate frame, which is centered at each agent’s position at the current timestep $t_{0}$.
 The transformed data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ has substantially reduced variance.
 We then standardize the transformed trajectories to zero mean and unit variance.
-To preserve the relative spatial relationships among agents, the reference states are collected as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$ consisting of each agent’s position at timestep $t_{0}$.
+To preserve the relative spatial relationships among agents, the reference states are collected as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$ consisting of each agent’s own position at timestep $t_{0}$.
 V-B Consistency Model Training
@@
 The other agents’ trajectories (always shown with ground truth) are in blue.
-Let $\mathcal{X}$ denote the space of transformed future trajectories $\Gamma(S_{e}^{\text{future}})$ and $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec.
+Let $\mathcal{X}$ denote the space of transformed future trajectories $\Gamma(S_{e}^{\text{future}})$ and $\Gamma(S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec.
 V-A.
@@
 (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec.
-IV-A as follows:
+IV-A with details in [38] as follows:
 $\mathcal{L}=\omega_{0}\mathcal{L}_{\text{consistency}}+\omega_{1}\mathcal{L}_{\text{encoder}}$ (3)
@@
 $c_{\text{goal}}=\sqrt{(p_{x}^{t_{0}+H_{2}}-p_{x,goal})^{2}+(p_{y}^{t_{0}+H_{2}}-p_{y,goal})^{2}}$
-$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
-$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (5)
+$c_{\text{acc}}=\frac{1}{H_{2}}\sum_{i=t_{0}+1}^{t_{0}+H_{2}}max(|a^{i}|-a_{\text{limit}},0)$
+$c_{\omega}=\frac{1}{H_{2}}\sum_{i=t_{0}+1}^{t_{0}+H_{2}}max(|\omega^{i}|-\omega_{\text{limit}},0)$ (5)
 $c_{\text{goal}}$ enforces goal-reaching for the ego vehicle at the final timestep $t_{0}+H_{2}$.
@@
 These constraints are inherently in conflict.
-For example, only moving the trajectory’s final state closer to the goal can increase acceleration violations of the car
+For example, only moving the trajectory’s final state closer to the goal can increase acceleration violations of the car.
 V-D Guided Sampling
@@
 It also simplifies step size tuning.
+Note that this differs from ADMM since it doesn’t have a dual update step.
 VI Experiments
@@
 TABLE I: Planning Trajectory Realism
-Method minADE ($m$) minFDE ($m$) Step
+Method minADE (m) minFDE (m) Sampling Step
+MotionDiffuser* [53] 0.533 0.007 32
 Transformer 1.237 2.796 1
@@
 We identify the surrounding agents with trajectories $S_{\mathcal{O}_{s}}$ by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold.
-This enables us to focus our predictive planner on the most relevant interactive behaviors between the ego vehicle and its immediate surroundings.
+This choice focuses the predictive planner on the most relevant and interactive neighboring agents, and our model architecture is fully compatible with more agents without additional modifications.
 The goal $s_{e}^{g}$ of the ego agent is chosen as its state at the end of the 8-second future horizon.
@@
 This baseline aims to accelerate the pretrained Diffusion-10 model with only 4 sampling steps in trajectory generation, with an efficient DDIM sampling strategy [41].
+MotionDiffuser[53]
+This is state-of-the-art diffusion-based approach for multi-agent trajectory prediction.
+As the official code has not been released, we report their results of trajectory realism directly from the original paper on the interactive split of the WOMD.
 Ours and Ours guided
@@
 Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints.
-In Table II, we evaluate the trajectory quality using four metrics 1) angle change as the smoothness of directional transitions, 2) path length as the efficiency of the chosen route, 3) curvature as the sharpness of turns, and 4) collision rate ($2m$ threshold) as safety.
+We also report MotionDiffuser [38] results on joint minADE and minFDE over the pair of interactive agents, which applies an attractor to the GT final point.
+Our method achieves comparable performance to MotionDiffuser while requiring significantly fewer sampling steps (4 vs.
+32).
+In Table II, we evaluate the trajectory quality using four metrics 1) Angle change, measuring the smoothness of directional transitions, 2) Path length, measuring the total distance traveled along the planning trajectory to reflect route efficiency, 3) Curvature, quantifying the sharpness of turns, and 4) Collision rate (with a $2m$ threshold), indicating safety.
 Lower values in these metrics indicate smoother, safer, and more efficient trajectories that align with natural driving patterns.
@@
 Our consistency model maintains high-quality trajectory planning when targeting novel goal locations not present in the dataset.
-The current transformer method (MTR) doesn’t support goal selections.
-With the same encoder, the computational time for diffusion-based methods directly depends on the number of sampling steps, shown in Table I.
-Diffusion-4 and DDIM have the same sampling step as our model, but fail to match our performance in trajectory generation across all metrics, as evidenced in Table I and Table II.
-The guided sampling, introducing a little computational overhead ($\approx 15\%$), provides higher quality trajectory generation and enables controllable planning.
-Even Diffusion-10, despite its more sampling steps, cannot achieve a comparable performance of trajectory generation to our approach in Table II.
+The current transformer method doesn’t support goal selections.
 VI-B3 Ablation Study on ADMM-inspired Guided Sampling
 We compare our ADMM-inspired method (ADMM for short) in Eq.
-(7) with the vanilla gradient descent (GD for short) in Eq.
+(7) with vanilla gradient descent (GD for short) in Eq.
 (6), which simultaneously optimizes all planning constraints.
-To illustrate the advantage of our method in hyperparameter tuning, we sweep through all combinations of step sizes ($\alpha=1e^{-4},1e^{-5},1e^{-6}$) across three constraints (27 combinations total), using 50, 100, and 150 gradient steps over 200 scenarios.
+To show the advantage of our method in hyperparameter tuning, we sweep through all combinations of step sizes ($\alpha=1e^{-4},1e^{-5},1e^{-6}$) across three constraints (27 combinations total), using 50, 100, and 150 gradient steps over 200 scenarios.
 TABLE III shows that GD results in significantly higher violations than ADMM on average; e.g., at 100 steps, GD has 9% higher goal-reaching, 35% higher acceleration, and 8% higher angular speed violations.
-The lowest sums of violations achieved are 1.45, 1.16, and 0.71 for 50, 100, and 150 steps, respectively.
-Increasing gradient steps reduces constraint violations but with computational tradeoffs.
+With increasing gradient steps of 50, 100, and 150 steps, we obtain lower sums of violation as 1.45, 1.16, and 0.71, but with more computational tradeoffs.
 In Fig.
@@
 In contrast, our ADMM method smoothly and effectively reduces all three violations.
+VI-B4 Computational Time
+With the same encoder, the computational time for diffusion-based methods directly depends on the number of sampling steps, shown in Table I.
+Diffusion-4 and DDIM have the same sampling step as our model, but fail to match our performance in trajectory generation across all metrics in Table I and Table II.
+The guided sampling, introducing a little computational overhead ($\approx 15\%$), provides higher quality trajectory generation and enables controllable planning.
+Even Diffusion-10, despite its more sampling steps, cannot achieve a comparable performance of trajectory generation to our approach in Table II.
 VII Conclusion